OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3?

OpenStudy (anonymous):

Please help soon.

OpenStudy (baru):

what can you say about the slope of parallel lines?

OpenStudy (anonymous):

they're equal, right?

OpenStudy (anonymous):

eq. of any line parallel to x+y=3 is x+y=a because it passes through (-1,-1) -1-1=a a=-2 reqd. eq. is x+y=-2